# A toric degeneration of Kronecker moduli spaces

### v1

@@ 4.1. Matching fields for Grassmannians
ranging over permutation $\sigma\in\operatorname{\mathrm{Sym}}(r)$, where $[x_{ij}]$ are the coordinates on an $r\times n$ matrix.
@@ 4.2. Generalizing to Kronecker moduli spaces
Fix a Kronecker quiver $\mathrm{K}^{n}_{r_{1}r_{2}}$, and a grading on $\mathbb{C}[x^{i}_{jk}]$ given by a tuple of matrices of non-negative integers $[c^{i}_{jk}]\in M$. Note that this grading gives a grading on the coordinates of $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$ and $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$. We’ll be interested in the situation when $[c^{i}_{jk}]$ induces matching fields simultaneously on $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$, $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$ and $\mathrm{K}^{n}_{r_{1}r_{2}}$, which we’ll denote $\Lambda_{1},$ $\Lambda_{2}$ and $\Lambda$ respectively.
@@
Let $[c^{i}_{jk}]\in M$ be a grading giving matching fields $\Lambda_{1},$ $\Lambda_{2}$ and $\Lambda$ on $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$, $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$, and $\mathrm{K}^{n}_{r_{1}r_{2}}$ respectively. Suppose the following two conditions hold:
@@
Note that defining canonical tableaux for Grassmannians implies a construction for Kronecker moduli. Given a coherent matching field for $\mathrm{K}^{n}_{r_{1}r_{2}}$, if we have defined canonical tableau for the associated coherent matching fields of $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$ and $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$ then we define a canonical linked pair $(T^{\pm})$ to be a linked pair such that both $T^{+}$ and $T^{-}$ are canonical. Then for each $v\in C_{1}\cap C_{2}$, there is a unique canonical linked pair $T^{\pm}$ such that $\mathrm{Mon}(T^{\pm})$ has exponent $v$. The other desired condition – that $\mathrm{LM}(f_{T^{\pm}})=\mathrm{Mon}(T^{\pm})$ for canonical pairs – is not obviously (to the author) true. Instead, we check it as part of the construction in examples.
@@
    Start by finding a grading $[c^{i}_{jk}]$ such that it induces matching fields on $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$ and $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$ for which the Plücker coordinates are a SAGBI basis in this order (for families where this is known, see [23, 4]).
@@
where $<$ is the lex ordering. Then the grading of the monomial corresponding to $\sigma\in\operatorname{\mathrm{Sym}}(r_{1})$ is
@@
where $<$ is the lex ordering. Then the grading of the monomial corresponding to $\sigma\in\operatorname{\mathrm{Sym}}(r_{2})$ is
@@ 5. Examples
$(\begin{bmatrix}1&2\\ 2&4\\ 3&6\\ \end{bmatrix},\hskip 8.53581pt\begin{bmatrix}4&8\\ 8&16\\ 12&24\\ \end{bmatrix},\hskip 8.53581pt\begin{bmatrix}16&32\\ 32&64\\ 48&96\\ \end{bmatrix}).$
@@
$(\begin{bmatrix}64&128\\ 128&256\\ 3&6\\ \end{bmatrix},\hskip 8.53581pt\begin{bmatrix}4&8\\ 8&16\\ 12&24\\ \end{bmatrix},\hskip 8.53581pt\begin{bmatrix}16&32\\ 32&64\\ 48&96\\ \end{bmatrix}).$

### v2

@@ 4.1. Matching fields for Grassmannians
ranging over permutation $\sigma\in\operatorname{\mathrm{Sym}}(r)$, where $x_{ij}]$ are the coordinates on an $r\times n$ matrix.
@@ 4.2. Generalizing to Kronecker moduli spaces
Fix a Kronecker quiver $\mathrm{K}^{n}_{r_{1}r_{2}}$, and a grading on $\mathbb{C}[x^{i}_{jk}]$ given by a tuple of matrices of non-negative integers $c^{i}_{jk}]\in M$. Note that this grading gives a grading on the coordinates of $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$ and $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$. We’ll be interested in the situation when $c^{i}_{jk}]$ induces matching fields simultaneously on $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$, $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$ and $\mathrm{K}^{n}_{r_{1}r_{2}}$, which we’ll denote $\Lambda_{1},$ $\Lambda_{2}$ and $\Lambda$ respectively.
@@
Let $c^{i}_{jk}]\in M$ be a grading giving matching fields $\Lambda_{1},$ $\Lambda_{2}$ and $\Lambda$ on $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$, $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$, and $\mathrm{K}^{n}_{r_{1}r_{2}}$ respectively. Suppose the following two conditions hold:
@@
Note that defining canonical tableaux for Grassmannians implies a construction for Kronecker moduli. Given a coherent matching field for $\mathrm{K}^{n}_{r_{1}r_{2}}$, if we have defined canonical tableaux for the associated coherent matching fields of $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$ and $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$ then we define a canonical linked pair $(T^{\pm})$ to be a linked pair such that both $T^{+}$ and $T^{-}$ are canonical. Then for each $v\in C_{1}\cap C_{2}$, there is a unique canonical linked pair $T^{\pm}$ such that $\mathrm{Mon}(T^{\pm})$ has exponent $v$. The other desired condition – that $\mathrm{LM}(f_{T^{\pm}})=\mathrm{Mon}(T^{\pm})$ for canonical pairs – is not obviously (to the author) true. Instead, we check it as part of the construction in examples.
@@
    Start by finding a grading $c^{i}_{jk}]$ such that it induces matching fields on $\operatorname{\mathrm{Gr}}(r_{1},nr_{2})$ and $\operatorname{\mathrm{Gr}}(r_{2},nr_{1})$ for which the Plücker coordinates are a SAGBI basis in this order (for families where this is known, see [23, 3]).
@@
where << is the lex ordering. Then the grading of the monomial corresponding to $\sigma\in\operatorname{\mathrm{Sym}}(r_{1})$ is
@@
where << is the lex ordering. Then the grading of the monomial corresponding to $\sigma\in\operatorname{\mathrm{Sym}}(r_{2})$ is
@@ 5. Examples
[…]
@@
[…]